OpenStudy (anonymous):

3x + 2y = 13 2x + 3y = 12 Determine the solution to the system and check your solution in both equations. Input the solution to the system, putting the solution in ordered pair (x, y) form.

OpenStudy (anonymous):

Multiply the first equation by 2 and the second -3 and ad them 6x + 4 y =26 -6x- 9y =-36 Add them -5y =-10 , y=2 replace in the first equation and find x.

OpenStudy (anonymous):

systems of equations

OpenStudy (anonymous):

replace in y the first equation and find x.

OpenStudy (anonymous):

EXACTLY

jhonyy9 (jhonyy9):

3x+2y=13 multiply by -2 2x+3y=12 ------- 3 -6x-4y=-26 6x+9y=36 -------------- assum 0 5y=10 y=10/5 =2 y=2 2x+3*2=12 2x=12-6 2x=6 x=6/2 x=3 x=3 y=2

jhonyy9 (jhonyy9):

(3,2) solutions pair